OpenStudy (anonymous):

What are all the real zeros of f(x)=x^4+2x^3+x^2-4

OpenStudy (anonymous):

factor it: hint\[f(x)=x^4+2x^3+x^2-4=x^4-1+2(x^3-1)+x^2-1\]